Lecture 2 - cont. Basic Probability

We start at [[handout01-ProbBasics-350F24.pdf#page=4]]:

(a) P(A′)=1−P(A)=.96
P(B′)=1−P(B)=.67

(b) A∩B are the set of submissions with both types of errors.

(c)

Use the diagram to get that it would be impossible to get P(A∩B) so far. Namely, we could have A totally within B:

This would make P(A∩B)=P(A)=.14. That would be the maximum, as A,B could be disjoint:

Then $P(A \cap B) = .0$ clearly.

Thus we have, with the current information that 0≤P(A∩B)≤.14.

(d) A∪B represents getting either a syntax error or a logic error (just any error for this case).

(e) P(A∪B)=P(A)+P(B)−P(A∩B) (since we don't want to count the intersection bit twice). As a result then since P(A∩B)∈[0,.14] then we have:

P(A∪B)=.14+.33−P(A∩B)=.47−P(A∩B)∈[.33,.47]

This is the same addition rule we derived separately.
(f) If we are given the P(A∩B)=.05 then we can use our formula to get:

P(A∪B)=.14+.33−.05=.47−.05=.42

(g) P(A′∩B′)=1−P(A∪B)=1−.42=.58.

(h) P(A′∩B)=P(A∪B)−P(A)=.42−.14=.28.

Contingency Table

We can make a contingency table for this situation:

Syntax Errors B No Syntax Errors B′
Logic Errors A .14
No Logic Errors A′
.33 1.0

With the supposition that P(A∩B)=.05:

Syntax Errors B No Syntax Errors B′
Logic Errors A .05 .09 .14
No Logic Errors A′ .28 .58 .86
.33 .67 1.0
The columns will add up to the bottom number, and the rows add to the last column.

(f) P(A∪B)=.14+.33−.05=.42 as we found before.
(g) P(A′∩B′)=.58
(h) P(A′∩B)=.28

A 3 Example

(a)


(b) The .07 probability seen in the diagram refers to A∩B′∩C′ which is a purely minor defect.